John and Jim are going on a bike ride. On the trip to mill park, they ride at an average of 8 mph on a route that is 4 miles less than the route on the return trip. They travel 7 mph on the return trip. If it takes them 1hour longer to return than to go, how much total time do they spend riding their bikes?
Answer by mananth(16949)

John and Jim are going on a bike ride. On the trip to mill park, they ride at an average of 8 mph on a route that is 4 miles less than the route on the return trip. They travel 7 mph on the return trip. If it takes them 1hour longer to return than to go, how much total time do they spend riding their bikes?
return trip distance be x miles
forward distance to mill park ---- x-4 miles
..
speed to mill park = 8 mph
return speed = 7 mph
..
return time -time to go to mill park = 1 hour
..
t=d/r
..
x/7-(x-4)/8 =1
(8x-7(x-4))/56=1
8x-7x+28=56
x=56-28
x=28 miles return distance
forward distance = 28-4 = 24 miles
..
t= d/r
28/7 + 24/8 = 7 hours
so They spend 7 hours riding
...
m.ananth@hotmail.ca